OpenStudy (anonymous):

LOOK IM HONESTLY DESPERATE. TODAY IS TODAY IS THE LAST DAY. PLEASE What is the slope of the line that passes through the points (3, 1) and (–2, 5)?

OpenStudy (mertsj):

Subtract the y values Subtract the x values Put the first number over the second number.

OpenStudy (anonymous):

quit procrastinating

OpenStudy (anonymous):

The slope is -5/4.

OpenStudy (anonymous):

thank you!!!!! Kevonni

OpenStudy (anonymous):

Select the equation of the line in slope-intercept form: 4x – 9y = –36

OpenStudy (mertsj):

The slope of the first one is NOT -5/4

OpenStudy (anonymous):

y = (4/9)x + 4 4x - 9y = -36 4x - 9y + 9y = -36 + 9y 4x = -36 + 9y 4x + 36 = -36 + 9y + 36 4x + 36 = 9y 9y = 4x + 36 9y/9 = 4x/9 + 36/9 y = (4/9)x + 4

OpenStudy (anonymous):

Thanks so much!!!! do you have time for more?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which of the following has a slope of –3 and a y-intercept of –5? a)y = –3x + 5 b) y = –3x – 5 c) y = 3x + 5 d) y = 3x – 5

OpenStudy (anonymous):

B

OpenStudy (anonymous):

The graph of a line has a slope of 2 and passes through the point (–6, –5). What is the equation of the line in point-slope form? a)y – 6 = 2(x – 5) b) y + 6 = 2(x + 5) c) y – 5 = 2(x – 6) d) y + 5 = 2(x + 6)